media%2F56b%2F56b0cd2c-3157-4a28-a478-ec
How much of the final amount is interest?anwer is part as well.
Starting at age 35, you deposit $2000 a year into an IRA account for retirement. Treat the yearly deposits into the account continuously, how much will be in the account 30 years later, when ou retire at age 65? How much of the final amount is What is the value of the IRA when you turn 65? S (Round to the nearest dollar as needed.)
an RA acountforreti e ment Treat the yearly deposits into the account as a continuous income stream years later, when you retire at age 65? How much of the final amount is interest? If money in the account earns 5%, compounded

How much of the final amount is interest?
Answer #1

Yearly installment amount : P = $2000

Rate of interest : r = 5% = 0.05

Number of payments made in 30 years : t = 30

Future value after t'' installments (if compounded continuously):

ert-1 FV = P *

Plug values into the formula:

.054301 FV 20005 135,814.89

Value of the IRA when you turn 65 is $135,814.89.

Total amount paid as installments in 30 years = 2000*30 = $60,000

Thus interest earned = FV - P = 135,814.89 - 60,000 = $75,817.89